OpenStudy (anonymous):

What is the limit of the sequence 9^(n+1)/10^n. Thanks :)

OpenStudy (aum):

\[\large \frac{9^{n+1}}{10^n} = \frac{9*9^{n}}{10^n} = 9*\left(\frac {9}{10}\right)^n \]

OpenStudy (aum):

As n->infinity, (9/10)^n -> ?

OpenStudy (anonymous):

0?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

Thank you!

OpenStudy (aum):

You are welcome.